OpenStudy (anonymous):

Which is not a solution to the equation -3x + y=5? A. (–4, –7) B. (3, 14) C. (2, –1) D. (1, 8)

OpenStudy (anonymous):

An ordered pair (x, y) will be a solution to the line if when you substitute the values for x and y into the equation, the left hand side will equal 5 in this case.

OpenStudy (anonymous):

Thought I told you this already o_O. So, try the first point, (-4,-7) -3(-3)+(-7) = 5...5=5 so yes this is your point. I guess you just needed an example.

OpenStudy (anonymous):

For example, A (-4, -7) is a solution since -3(-4) + (-7) = 12 - 7 = 5

OpenStudy (anonymous):

eehh no that was a different one this is a new one oh and thanks for the help on the last one tho

OpenStudy (anonymous):

thanks gryphon

OpenStudy (anonymous):

You are welcome!